If x = 3 is the solution of x^2 + kx +15, value of k is
1
Expert's answer
2020-08-17T18:29:44-0400

"x^2+kx+15=0"

If "x=3" is the solution of this equation, then "3^2+3\times k+15=24+3k=0" and "k=-8."


Answer: "k=-8."